OpenStudy (gabylovesyou):

Which of the following is equivalent to 48xy2 + 24xy4 − 12x2y4? 12xy2(4 + 2y2 − xy2) 12xy2(4 − 2y2 + xy2) 12xy2(3 + 2y2 − xy2) 12xy2(4 − 3y2 − xy2)

OpenStudy (gabylovesyou):

@jim_thompson5910 @ganeshie8

OpenStudy (zzr0ck3r):

what can you factor out of your equation? 12xy^2? then what do you have left?

OpenStudy (luigi0210):

Well all the problems seem to factor a 12xy^2, try doing that and see what you get

OpenStudy (dan815):

what would shakira do

OpenStudy (dan815):

always ask yourself this question

jimthompson5910 (jim_thompson5910):

how can you factor 48 where one of the factors is 12

jimthompson5910 (jim_thompson5910):

in other words, what is the solution to the equation 48 = 12x

OpenStudy (anonymous):

it is a just take 12xy2 common from it

OpenStudy (gabylovesyou):

x = 4

jimthompson5910 (jim_thompson5910):

so 48 = 12*4

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

this means that if you had something like 48x + 12y, you can write it like this 48x + 12y 12*4x + 12*1y 12(4x + 1y) ... use the distributive property to factor out 12 12(4x + y)

jimthompson5910 (jim_thompson5910):

or if you had something like this 48x + 36y, you can do this 48x + 36y 12*4x + 12*3y 12(4x + 3y)

OpenStudy (gabylovesyou):

ok :)

jimthompson5910 (jim_thompson5910):

tell me what you get when you factor the original expression up at the top

OpenStudy (gabylovesyou):

factor as in foil or ...

jimthompson5910 (jim_thompson5910):

factor as in use the distributive property

OpenStudy (gabylovesyou):

48x + 2y

jimthompson5910 (jim_thompson5910):

you're on the right track, but you are a bit off

OpenStudy (gabylovesyou):

how ?

jimthompson5910 (jim_thompson5910):

you have 48xy^2 how can you factor this so 12xy^2 is one factor

OpenStudy (gabylovesyou):

ummm idk ;/

jimthompson5910 (jim_thompson5910):

well 48 factors to 12*4 right?

OpenStudy (gabylovesyou):

yes ..... and 6 * 2 factors to 12 (if u needed that) lol

jimthompson5910 (jim_thompson5910):

yes we'll need that as well

jimthompson5910 (jim_thompson5910):

but basically what I'm trying to get at is that 48xy^2 factors to 4*12xy^2 = (12xy^2)*4

OpenStudy (gabylovesyou):

ohhhhhhh ok !!!!!! i guess i need the baby words like how many times does 12 go into 48 lol xD

jimthompson5910 (jim_thompson5910):

that's ok, approach it the best way for you

jimthompson5910 (jim_thompson5910):

how would we factor 24xy^4 where 12xy^2 is one factor

OpenStudy (gabylovesyou):

dividing ? hold on what r u trying to ask me ? :p

jimthompson5910 (jim_thompson5910):

yes in a way you will divide basically I want you to solve for p in the equation 24xy^4 = p*12xy^2

OpenStudy (gabylovesyou):

what times 12 is 24 ? thats what ur trying to say ?

jimthompson5910 (jim_thompson5910):

yes along with the question "what times xy^2 gives me xy^4"

OpenStudy (gabylovesyou):

4y^2 ?

jimthompson5910 (jim_thompson5910):

very good 4y^2 times 12xy^2 gives 48xy^4

jimthompson5910 (jim_thompson5910):

oops I was looking at the wrong piece

jimthompson5910 (jim_thompson5910):

it should be 2y^2 so 24xy^4 = 2y^2*12xy^2

jimthompson5910 (jim_thompson5910):

see how I'm getting this?

OpenStudy (gabylovesyou):

ohh idk where i got 4 from if 12 * 2 = 24 lol yes

jimthompson5910 (jim_thompson5910):

finally, 12x^2y^4 = q*12xy^2 what goes in place of q?

OpenStudy (gabylovesyou):

xy^2 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you're getting the hang of this

OpenStudy (gabylovesyou):

yaay

jimthompson5910 (jim_thompson5910):

48xy^2 + 24xy^4 − 12x^2y^4 turns into 4*12xy^2 + 2y^2*12xy^2 - xy^2*12xy^2 then you factor out the common term 12xy^2 to get what?

OpenStudy (gabylovesyou):

4 + 2y2 - xy^2

jimthompson5910 (jim_thompson5910):

you got it, you basically erase all the copies of 12xy^2 to get 4+2y^2-xy^2 then you write all that in a set of parenthesis with 12xy^2 sitting out front like this 12xy^2( 4+2y^2-xy^2)